# A Bayesian approach to functional regression: theory and computation

### v2

@@ 2 A Bayesian methodology for RKHS-based functional regression models
The functional models contemplated here are those obtained by considering a functional parameter $\alpha\in\mathcal{H}(K)$ and replacing the scalar product for $\left\langle X,\alpha\right\rangle_{K}$ in the $L^{2}$-models (1.1) and (1.2). As previously explained, this change in perspective has tangible benefits both in theory and practice on account of the RKHS properties. However, to further simplify things we will follow a parametric approach and suppose that $\alpha$ is in fact a member of the dense subspace $\mathcal{H}_{0}(K)$ defined in (1.3). Although natural in this context, this crucial assumption may seem too restrictive at first glance. Nonetheless, we will demonstrate that this scheme has a firm theoretical backing and that the resulting models perform sufficiently well in applied settings, seeming not to need the comprehensiveness (and complications) of the larger $L^{2}$-spaces.
As we said before, with a slight abuse of notation we will understand the expression $\left\langle x,\alpha\right\rangle_{K}$ as $\Psi_{x}(\alpha)$, where $x=X(\omega)$ and $\Psi_{x}$ is Loève’s isometry. Hence, taking into account that $\alpha\in\mathcal{H}_{0}(K)$ and that $\Psi_{X}(K(t,\cdot))=X(t)$ by definition, we can write $\left\langle x,\alpha\right\rangle_{K}\equiv\sum_{j}\beta_{j}x(t_{j})$ when $\alpha(\cdot)=\sum_{j}\beta_{j}K(t_{j},\cdot)$ and $x$ is a realization of $X$. In this way we get a simpler, finite-dimensional approximation of the functional RKHS model, which we argue reduces the overall complexity while still capturing most of the relevant information. By restricting the ambient space to $\mathcal{H}_{0}(K)$ we are essentially truncating the value of $p$, the dimensionality, and thus advocating for parsimony. Moreover, the model remains “truly functional” in the sense that $p$ is variable: we are exploiting the RKHS perspective to give a functional nature to all finite-dimensional models based on linear combinations of the marginals. Even though we are selecting a set of impact points $\{t_{j}\}$ that depends on the observed grid, we can circumvent problems with this approach by smoothing or interpolating the functional data in a pre-processing step.
@@
In view of (1.3) and Loève’s isometry, to set a prior distribution on the unknown function $\alpha$ (that is, a prior distribution on the functional space $\mathcal{H}_{0}(K)$) it suffices to consider a discrete distribution on the number of components $p$, and then impose $p$-dimensional continuous prior distributions on the coefficients $\beta_{j}$ and the times $t_{j}$ given $p$. Thanks to this parametric approach, the challenging task of setting a prior distribution on a space of functions is considerably simplified, while simultaneously not constraining the model to any specific distribution (in contrast to, say, Gaussian process regression methods). Moreover, note that our simplifying assumption on $\alpha$ is not actually very strong, since any prior distribution $\mathbb{P}_{0}$ on $\mathcal{H}_{0}(K)$ can be directly extended to a prior distribution $\mathbb{P}$ on $\mathcal{H}(K)$: just define $\mathbb{P}(B)=\mathbb{P}_{0}(B\cap\mathcal{H}_{0}(K))$ for all Borel sets $B$ on $\mathcal{H}(K)$.
@@
Lastly, since the value of $p$ can vary, we need a way to introduce dimension information in our MCMC posterior approximation scheme. There are several alternatives in the literature, such as product space formulations (Carlin and Chib, 1995) or Bayesian averaging/model selection methods (Hoeting et al., 1999), but this is precisely the problem that reversible jump samplers were designed to solve. The use of these samplers fits naturally within our framework, as they allow the complexity of the model to be directly chosen by the data, jointly inferring about the dimensionality and the value of the parameters. We explore this approach in some detail in Section 2.3.
@@ 2.1 Functional linear regression
$Y=\alpha_{0}+\left\langle X,\alpha\right\rangle_{K}+\varepsilon=\alpha_{0}+\sum_{j=1}^{p}\beta_{j}X(t_{j})+\varepsilon,$ (2.1)
where $\alpha(\cdot)=\sum_{j=1}^{p}\beta_{j}K(t_{j},\cdot)\in\mathcal{H}_{0}(K)$, $\alpha_{0}\in\mathbb{R}$, and $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ is an error term independent from $X$. This model is essentially a finite-dimensional approximation from a functional perspective to the more general RKHS model that assumes $\alpha\in\mathcal{H}(K)$ (e.g. Berrendero et al., 2024). Since the number of components $p$ is unknown, the full parameter space is $\Theta=\bigcup_{p\in\mathbb{N}}\Theta_{p}$, where $\Theta_{p}=\mathbb{R}^{p}\times[0,1]^{p}\times\mathbb{R}\times\mathbb{R}^{+}$. In the sequel, a generic element of $\Theta_{p}$ will be denoted by $\theta_{p}=(\beta_{1},\dots,\beta_{p},t_{1},\dots,t_{p},\alpha_{0},\sigma^{2})\equiv(b_{p},\tau_{p},\alpha_{0},\sigma^{2})$, though we will occasionally omit the subscript when the value of $p$ is understood. Before proceeding further, observe that we can rewrite model (2.1) in a more explicit and practical fashion in terms of the available sample information. For $\theta\in\Theta$, the reinterpreted model assumes the form
$Y_{i}\mid X_{i},\theta\ \stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\mathcal{N}\left(\alpha_{0}+\sum_{j=1}^{p}\beta_{j}X_{i}(t_{j}),\ \sigma^{2}\right),\quad i=1,\dots,n.$ (2.2)
It is worth mentioning that the model remains linear in the sense that it fundamentally involves a random variable $\left\langle X,\alpha\right\rangle_{K}=\Psi_{X}(\alpha)$ belonging to the linear span of the process $X$ in $L^{2}(\Omega)$. Also, note that given the number of components $p$ and the time instants $t_{j}$, the model becomes a multiple linear model with the $X(t_{j})$ as scalar covariates. In fact, this RKHS model is particularly suited as a basis for variable selection methods (Berrendero et al., 2019), and furthermore the general RKHS model entails the classical $L^{2}$-model (1.1) under certain conditions (see Berrendero et al., 2024). In addition, this model could be easily extended to the case of several covariates via an expression of type $Y=\alpha_{0}+\Psi_{X^{1}}(\alpha_{1})+\cdots+\Psi_{X^{q}}(\alpha_{q})+\varepsilon$. Then, as argued in Grollemund et al. (2019) for a similar situation, we could recover the full posterior by looking alternately at the posterior distribution of each covariate conditional on the rest of them.
@@
A simple and intuitive prior distribution for the parameter vector $\theta\in\Theta$, suggested by the structure of the parameter space and usually employed in the literature, is given by
$p$ $\sim\pi(p),$ (2.3)
$t_{j}$ $\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\mathcal{U}[0,1],\quad$ $j$ $=1,\dots,p,$
$\beta_{j}$ $\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\mathcal{N}(0,\eta^{2}),\quad$ $j$ $=1,\dots,p,$
$\pi(\alpha_{0},\sigma^{2})$ $\propto 1/\sigma^{2},$
where $\pi(p)$ is any discrete distribution on $\mathbb{N}$ (e.g. uniform on a given finite set) and $\eta^{2}\in\mathbb{R}^{+}$ is a hyperparameter of the model that depends strongly on the expected scale of the data. On the one hand, note the use of a joint prior distribution on $\alpha_{0}$ and $\sigma^{2}$, which is a widely used non-informative prior known as Jeffrey’s prior (Jeffreys, 1946). One should be wary of the Jeffreys-Lindley paradox when assigning improper priors (see Robert, 2014), but since all the parameters involved are common to all possible values of $p$, this formulation should not present difficulties. On the other hand, the prior on the coefficients $\beta_{j}$ is deliberately kept simple and somewhat non-informative, as it will greatly speed up and simplify the computations of the RJMCMC sampler later on.
@@ 2.2 Functional logistic regression
In the case of functional logistic regression, we regard the binary response variable $Y\in\{0,1\}$ as a Bernoulli random variable given the regressor $X=x\in L^{2}[0,1]$, and as usual suppose that $\log\left(p(x)/(1-p(x))\right)$ is linear in $x$, where $p(x)=\mathbb{P}(Y=1|X=x)$. Then, following the approach suggested by Berrendero et al. (2023), a simplified logistic RKHS model is given, in terms of the correspondence $\left\langle X,\alpha\right\rangle_{K}=\Psi_{X}(\alpha)$, by the equation
$\mathbb{P}(Y=1\mid X)=\frac{1}{1+\exp\{-\alpha_{0}-\left\langle X,\alpha\right\rangle_{K}\}},\quad\alpha_{0}\in\mathbb{R},\ \alpha\in\mathcal{H}_{0}(K).$ (2.4)
Indeed, note that this can be seen as a finite-dimensional approximation (with a functional interpretation) to the general RKHS functional logistic model proposed by these authors, which can be obtained by replacing $\mathcal{H}_{0}(K)$ with $\mathcal{H}(K)$. Now, if we aim at a classification problem, our strategy will be similar to that followed in the functional linear model: after incorporating the sample information, we can rewrite (2.4) as
$Y_{i}\mid X_{i},\theta\ \stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\operatorname{Bernoulli}(p_{i}),\quad i=1,\dots,n,$ (2.5)
with
$p_{i}=\mathbb{P}(Y_{i}=1\mid X_{i},\theta)=\frac{1}{\displaystyle 1+\exp\left\{-\alpha_{0}-\sum_{j=1}^{p}\beta_{j}X_{i}(t_{j})\right\}},\quad i=1,\dots,n,$
where in turn $\alpha_{0},\beta_{j}\in\mathbb{R}$ and $t_{j}\in[0,1]$.
@@
As far as prior distributions go, we proceed as we did in the linear model. However, following the advice in Gelman et al. (2008) and Ghosh et al. (2018) we do change the prior of the coefficients $\beta_{j}$ and the constant $\alpha_{0}$, since their interpretation is different now:
$p$ $\sim\pi(p),$ (2.6)
$t_{j}$ $\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\mathcal{U}[0,1],\quad$ $j$ $=1,\dots,p,$
$\beta_{j}$ $\stackrel{{\scriptstyle\text{ind.}}}{{\sim}}t_{5}(0,2.5),\quad$ $j$ $=1,\dots,p,$
$\alpha_{0}$ $\sim\text{Cauchy}(0,10).$
The scaled Student’s $t$ and Cauchy distributions represent a robust and weakly informative prior that provides shrinkage and can handle the case of separation in logistic regression. The specific values of the parameters have been chosen via experimentation following the initial recommendations in Ghosh et al. (2018), which also establishes the need to scale the regressors to have mean 0 and standard deviation 0.5. Lastly, note that in this case the nuisance parameter $\sigma^{2}$ does not appear.
@@ 2.3 Reversible jump samplers for prediction
$\alpha_{\text{nested}}=\min\left(1,\frac{\mathcal{L}(\theta_{p+1})}{\mathcal{L}(\theta_{p})}\frac{\pi(\theta_{p+1})}{\pi(\theta_{p})}\frac{b_{p,p+1}}{d_{p+1,p}}\frac{1}{q(\theta_{+1})}\right).$
Here $p$ is the current number of components, $b_{p,p+1}$ is the probability of increasing the dimension from $p$ to $p+1$, $d_{p+1,p}$ is the probability of the reverse move (death), $q(\theta_{+1})$ is the proposal distribution for the added source, $\pi$ is the prior probability and $\mathcal{L}$ represents the likelihood. This is already a rather manageable expression that can be implemented efficiently, but we introduce another simplification: the proposal distribution is chosen to match the prior distribution, so that the corresponding terms cancel out. Nonetheless, in real-world scenarios one could consider more sophisticated proposal distributions that might better explore the parameter space; see for example Davies et al. (2023) or Korsakova et al. (2024) for some interesting ideas.
From a higher level perspective, the RJMCMC algorithm is an iterative procedure that produces a chain of $M$ approximate samples $(p_{m},\theta^{*}_{p_{m}})$ of the posterior distribution $\pi(p,\theta_{p}|\mathcal{D}_{n})$, each possibly of a different dimension; see Appendix C.3 for a visual representation. Given a previously unseen regressor $X_{\text{test}}$, with each of these samples we can generate an individual response following our RKHS models, denoted by $F(X_{\text{test}},\theta^{*}_{p_{m}})$. In the linear case we sample from $Y|X_{\text{test}},\theta^{*}_{p_{m}}$ as in (2.2), and in the logistic case we directly consider the probabilities $\mathbb{P}(Y|X_{\text{test}},\theta^{*}_{p_{m}})$ in (2.5) instead. We propose to combine these predictions in four different ways.
@@
The first idea involves utilizing all available information by summarizing and aggregating every individual RKHS response. We start by considering a point summary statistic $g$ (e.g. mean, median, mode) and consolidating the predictions from all sub-models, each having distinct dimension. Then, we bring together these predictions through a weighted sum, in which the weights correspond to the relative frequencies of the corresponding values of $p$:
$\hat{Y}=\sum_{p}\tilde{\pi}(p\mid\mathcal{D}_{n})g(\{F(X_{\text{test}},\theta^{*}_{p_{m}})\}_{m:p_{m}=p}),$
where $\tilde{\pi}(p|\mathcal{D}_{n})=M^{-1}\sum_{m}\mathbb{I}(p_{m}=p)$ and $\mathbb{I}$ is the indicator function. Note that this approach is reminiscent of the factorization of the posterior distribution as $\pi(p,\theta_{p}|\mathcal{D}_{n})=\sum_{p}\pi(p|\mathcal{D}_{n})\pi(\theta_{p}|p,\mathcal{D}_{n})$. Additionally, in the logistic case we employ the usual thresholding procedure to convert the final probability to a binary class label in $\{0,1\}$. We shall see in Section 4 that this method produces the best results in practice.
@@
We also consider a MAP strategy, where only the most probable sub-model is used and the rest of the samples are discarded. In this case, if $\tilde{p}=\arg\max_{p}\tilde{\pi}(p|\mathcal{D}_{n})$, predictions are computed as
$\hat{Y}=g(\{F(X_{\text{test}},\theta^{*}_{p_{m}})\}_{m:p_{m}=\tilde{p}}).$
Although this method potentially ignores many samples in the posterior approximation, this omission may help reduce the noise and prevent outliers from affecting the final prediction.
@@
In these methods we focus only on the marginal posterior distribution of $\tau_{p}=(t_{1},\dots,t_{p})$, disregarding the rest of the parameters and effectively constructing a variable selection procedure. After choosing a set of time instants $\{\hat{t}_{j}\}$ using the approximate posterior samples $\{\tau^{*}_{p_{m}}\}$, we can reduce the original regressors to just the marginals $\{X_{i}(\hat{t}_{j})\}$ and apply any of the well-known finite-dimensional prediction algorithms suited for this situation. We have again two different strategies.
@@
$\hat{Y}=\sum_{p}\tilde{\pi}(p\mid\mathcal{D}_{n})H(X_{\text{test}},\hat{\tau}_{p}),$
where $\hat{\tau}_{p}=(\hat{t}_{1},\dots,\hat{t}_{p})=g(\{\tau^{*}_{p_{m}}\}_{m:p_{m}=p})$ and $g$ is a component-wise summary statistic. The function $H$ represents the prediction for $X_{\text{test}}$ of a regular linear/logistic regression algorithm, fitted with the transformed data set $\{(X_{i}(\hat{\tau}_{p}),Y_{i}):i=1,\dots,n\}$. As before, in the logistic case the responses are thresholded to obtain a class label.
@@
We also consider a MAP approach to variable selection with only the information of the most probable model, i.e.,
$\hat{Y}=H(X_{\text{test}},\hat{\tau}_{\tilde{p}}).$

### v3

@@ 2 A Bayesian methodology for RKHS-based functional regression models
In principle, the functional RKHS models contemplated in this work are those obtained by considering a functional parameter $\beta\in\mathcal{H}(K)$ and replacing the scalar product for $\left\langle X,\beta\right\rangle_{K}$ in the $L^{2}$-models (1.1) and (1.2), which has tangible benefits both in theory and practice on account of the RKHS properties. However, to further simplify things we will follow a parametric approach and suppose that $\beta$ is in fact a member of the dense subspace $\mathcal{H}_{0}(K)$ defined in (1.3), i.e.:
$\beta(\cdot)=\sum_{j=1}^{p}\beta_{j}K(t_{j},\cdot).$ (2.1)
As we said before, with a slight abuse of notation we will understand the expression $\left\langle x,\beta\right\rangle_{K}$ as $\Psi_{x}(\beta)$, where $x=X(\omega)$ is a realization of $X$ and $\Psi_{x}$ is Loève’s isometry. Hence, taking into account that $\beta(\cdot)=\sum_{j}\beta_{j}K(t_{j},\cdot)$ and that $\Psi_{X}(K(t,\cdot))=X(t)$ by definition, we can identify $\left\langle x,\beta\right\rangle_{K}$ with $\sum_{j}\beta_{j}x(t_{j})$. In addition, we will assume that $X$ has a version with continuous sample paths, so that point evaluation is well-defined.
Although natural in this context, the crucial assumption that $\beta\in\mathcal{H}_{0}(K)$ may seem too restrictive at first glance, since we are essentially truncating the dimensionality of the model. Nevertheless, in this way we get a simpler, finite-dimensional approximation of the functional RKHS model, which we argue reduces the overall complexity while still capturing most of the relevant information. Plus, the simplified model remains “truly functional” in the sense that the number of components $p$ and the time instants $t_{j}$ are not fixed beforehand. In short, we are exploiting the RKHS perspective to give a functional nature to the collection of finite-dimensional models based on random linear combinations of the marginals, offering a unified treatment of all the so-called impact points models (Lindquist and McKeague, 2009; Kneip et al., 2016; Poß et al., 2020, see).
@@
In view of (2.1), to place a prior distribution on the unknown function $\beta$ (that is, a prior distribution on the functional space $\mathcal{H}_{0}(K)$) it suffices to consider a discrete distribution on the number of components $p$, and then select $p$-dimensional continuous priors for the coefficients $\beta_{j}$ and the times $t_{j}$ given $p$. Thanks to this parametric approach, the challenging task of setting a prior distribution on a space of functions is considerably simplified, while simultaneously not constraining the model to any specific distribution (in contrast to, say, Gaussian process regression methods). Moreover, note that our simplifying assumption on $\beta$ is not particularly limiting from a Bayesian point of view, since any distribution $\mathbb{P}_{0}$ on $\mathcal{H}_{0}(K)$ can be directly extended to a distribution $\mathbb{P}$ on $\mathcal{H}(K)$ by defining $\mathbb{P}(B)=\mathbb{P}_{0}(B\cap\mathcal{H}_{0}(K))$ for all Borel sets $B$ on $\mathcal{H}(K)$.
@@
Lastly, since the value of $p$ can vary, we need a way to introduce dimension information in our MCMC posterior approximation scheme. There are several alternatives such as product space formulations (Carlin and Chib, 1995) or Bayesian averaging/model selection methods (Hoeting et al., 1999), but this is precisely the problem that reversible jump samplers were designed to solve. The use of these samplers fits naturally within our framework, as they allow the complexity of the model to be directly chosen by the data, jointly inferring about the dimensionality and the value of the parameters.
@@ 2.1 Functional linear regression
$Y=\alpha+\left\langle X,\beta\right\rangle_{K}+\varepsilon=\alpha+\sum_{j=1}^{p}\beta_{j}X(t_{j})+\varepsilon,$
where $\beta(\cdot)=\sum_{j=1}^{p}\beta_{j}K(t_{j},\cdot)\in\mathcal{H}_{0}(K)$, $\alpha\in\mathbb{R}$, and $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$ is an error term independent from $X$. This model is essentially a finite-dimensional approximation from a functional perspective to the more general RKHS model that assumes $\beta\in\mathcal{H}(K)$ (Berrendero et al., 2024). Since the number of components $p$ is unknown, the full parameter space is $\Theta=\bigcup_{p\in\mathbb{N}}\Theta_{p}$, where $\Theta_{p}=\mathbb{R}^{p}\times[0,1]^{p}\times\mathbb{R}\times\mathbb{R}^{+}_{0}$. In the sequel, a generic element of $\Theta_{p}$ will be denoted by $\theta_{p}=(\beta_{1},\dots,\beta_{p},t_{1},\dots,t_{p},\alpha,\sigma^{2})\equiv(b_{p},\tau_{p},\alpha,\sigma^{2})$, though we will occasionally omit the subscript when the value of $p$ is understood. For $\theta\in\Theta$, the reinterpreted model in terms of the available sample information is
$Y_{i}\mid X_{i},\theta\ \stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\mathcal{N}\left(\alpha+\sum_{j=1}^{p}\beta_{j}X_{i}(t_{j}),\,\sigma^{2}\right),\quad i=1,\dots,n.$ (2.2)
It is worth mentioning that the model remains linear in the sense that it fundamentally involves a random variable $\left\langle X,\beta\right\rangle_{K}=\Psi_{X}(\beta)$ in the linear span of the process $X$. Moreover, this RKHS model is particularly suited as a basis for variable selection methods (see Berrendero et al., 2019).
@@
A simple and intuitive prior distribution $\Pi$ for the parameter vector $\theta\in\Theta$, suggested by the structure of the parameter space and usually employed in the literature, is given by
[…]
where $\pi$ is any discrete distribution on $\mathbb{N}$ (e.g. uniform on a given finite subset) and $\eta^{2}\in\mathbb{R}^{+}$ is a hyperparameter of the model that depends strongly on the expected scale of the data. On the one hand, note the use of a joint prior distribution on $\alpha$ and $\sigma^{2}$, which is a widely used non-informative prior known as Jeffrey’s prior (Jeffreys, 1946). One should be wary of the Jeffreys-Lindley paradox when assigning improper priors (see Robert, 2014), but since the two parameters involved are common to all possible values of $p$, this formulation should not present difficulties. On the other hand, the prior on the coefficients $\beta_{j}$ (which are interchangeable when coupled with their respective $t_{j}$) is deliberately kept simple and independent of $p$, as this will greatly speed up and simplify the computations of the RJMCMC sampler later on. Nevertheless, the methods could be adapted to more complicated and dependent priors without too much trouble.
@@ 2.2 Functional logistic regression
In the case of functional logistic regression, we regard the binary response variable $Y\in\{0,1\}$ as a Bernoulli random variable given the functional regressor $X$. Then, following the approach suggested by Berrendero et al. (2023), a simplified logistic RKHS model is given by the equation
$\mathbb{P}(Y=1\mid X)=\frac{1}{1+\exp\{-\alpha-\left\langle X,\beta\right\rangle_{K}\}},\quad\alpha\in\mathbb{R},\ \beta\in\mathcal{H}_{0}(K).$ (2.4)
Indeed, note that this can be seen as a finite-dimensional approximation (with a functional interpretation) to the general RKHS functional logistic model proposed by these authors, which can be obtained by replacing $\mathcal{H}_{0}(K)$ with $\mathcal{H}(K)$. After incorporating the sample information, we can rewrite (2.4) in parametric form for $\theta\in\Theta$ as
$Y_{i}\mid X_{i},\theta\ \stackrel{{\scriptstyle\text{ind.}}}{{\sim}}\operatorname{Bernoulli}(H_{\theta}(X_{i})),\quad i=1,\dots,n,$ (2.5)
where
$H_{\theta}(X_{i})=\mathbb{P}(Y_{i}=1\mid X_{i},\theta)=\frac{1}{\displaystyle 1+\exp\left\{-\alpha-\sum_{j=1}^{p}\beta_{j}X_{i}(t_{j})\right\}},\quad i=1,\dots,n.$
@@
As far as prior distributions go, we proceed as we did in the linear model. However, following the advice in Gelman et al. (2008) and Ghosh et al. (2018), we do change the prior of the coefficients $\beta_{j}$ and the constant term $\alpha$, since their interpretation is different now:
[…]
The scaled Student’s $t$ and Cauchy distributions represent robust and weakly informative priors that provide shrinkage and can handle the case of separation in logistic regression. The specific values of the parameters have been chosen via experimentation following the initial recommendations in Ghosh et al. (2018), which also establishes the need to scale the regressors to have mean 0 and standard deviation 0.5. Note that in this case the nuisance parameter $\sigma^{2}$ does not appear in the model.
@@ 2.3 Reversible jump samplers for prediction
$\alpha_{\text{nested}}=\min\left(1,\frac{\mathcal{L}(Y|X,\theta_{p+1})}{\mathcal{L}(Y|X,\theta_{p})}\frac{\Pi(\theta_{p+1})}{\Pi(\theta_{p})}\frac{b_{p,p+1}}{d_{p+1,p}}\frac{1}{q(\theta_{+1})}\right).$
Here $p$ is the current number of components, $b_{p,p+1}$ is the probability of increasing the dimension from $p$ to $p+1$ (birth), $d_{p+1,p}$ is the probability of the reverse move (death), $q(\theta_{+1})$ is the proposal distribution for the added source, $\Pi$ is the prior probability and $\mathcal{L}$ represents the likelihood. This is already a rather manageable expression that can be implemented efficiently, but we introduce another simplification: the proposal distribution is chosen to match the prior distribution, so that the corresponding terms cancel out. Nonetheless, in real-world scenarios one could consider more sophisticated proposal distributions that might better explore the parameter space; see for example Davies et al. (2023) or Korsakova et al. (2024) for some interesting ideas.
From a higher level perspective, the RJMCMC algorithm is an iterative procedure that produces a chain of $M$ approximate samples $(p_{m},\theta^{*}_{p_{m}})$ of the posterior distribution $\Pi_{n}(p,\theta_{p}|\mathcal{D}_{n})$, each possibly of a different dimension; see Appendix C.3 for a visual representation. Given a previously unseen regressor $X_{\text{test}}$, with each of these samples we can generate an individual approximate response following our RKHS models, denoted by $F(X_{\text{test}},\theta^{*}_{p_{m}})$. In the linear case we sample from $Y|X_{\text{test}},\theta^{*}_{p_{m}}$ as in (2.2), and in the logistic case we directly consider the probabilities $\mathbb{P}(Y=1|X_{\text{test}},\theta^{*}_{p_{m}})$ in (2.5). We propose to combine these predictions in four different ways.
@@
The first idea involves utilizing all available information by summarizing and aggregating every individual RKHS response. We start by considering a point summary statistic $g$ (such as the mean, median or mode) and consolidating the predictions from all sub-models, each having distinct dimension. Then, we bring together these predictions through a weighted sum, in which the weights correspond to the relative frequencies of the corresponding values of $p$ (their approximate posterior):
$\hat{Y}=\sum_{p}\tilde{\Pi}_{n}(p\mid\mathcal{D}_{n})g(\{F(X_{\text{test}},\theta^{*}_{p_{m}})\}_{m:p_{m}=p}),$
where $\tilde{\Pi}_{n}(p|\mathcal{D}_{n})=M^{-1}\sum_{m}\mathbb{I}(p_{m}=p)$ and $\mathbb{I}$ is the indicator function. In the logistic case, we employ the usual thresholding procedure to convert the final probability to a binary class label in $\{0,1\}$. Note that this approach is reminiscent of the factorization $\Pi_{n}(p,\theta_{p}|\mathcal{D}_{n})=\sum_{p}\Pi_{n}(p|\mathcal{D}_{n})\Pi_{n}(\theta_{p}|p,\mathcal{D}_{n})$ of the posterior distribution. We shall see in Section 4 that this method produces the best results in practice.
@@
[…]
@@
In these methods we focus only on the marginal posterior distribution of $\tau_{p}=(t_{1},\dots,t_{p})$, disregarding the rest of the parameters and effectively constructing a variable selection procedure. After choosing a set of time instants $\hat{\tau}_{p}=(\hat{t}_{1},\dots,\hat{t}_{p})$ using the approximate posterior samples $\{\tau^{*}_{p_{m}}\}$, we can reduce the original functional regressors to just the marginals $\{X_{i}(\hat{\tau}_{p})\}=\{(X_{i}(\hat{t}_{1}),\dots,X_{i}(\hat{t}_{p}))\}$ and apply any of the well-known finite-dimensional prediction algorithms suited for this situation.
@@
$\hat{Y}=\sum_{p}\tilde{\Pi}_{n}(p\mid\mathcal{D}_{n})G(X_{\text{test}},\hat{\tau}_{p}),$
where $\hat{\tau}_{p}=(\hat{t}_{1},\dots,\hat{t}_{p})=g(\{\tau^{*}_{p_{m}}\}_{m:p_{m}=p})$ and $g$ is a component-wise summary statistic. The function $G$ represents the prediction for $X_{\text{test}}$ of a regular linear/logistic regression algorithm, fitted with the transformed data set $\{(X_{i}(\hat{\tau}_{p}),Y_{i}):i=1,\dots,n\}$.
@@
[…]